You are the manager of an organization in America that distributes blood to hospitals in all 50 states and the District of Columbia. A recent report indicates that nearly 50 Americans contract HIV each year through blood transfusions. Although every pint of blood donated in the United States undergoes a battery of nine different tests, existing screening methods can detect only the antibodies produced by the body’s immune system— not foreign agents in the blood. Since it takes weeks or even months for these antibodies to build up in the blood, newly effected HIV donors can pass along the virus through blood that has passed existing screening tests. Happily, researchers have developed a series of new tests aimed at detecting and removing infections from donated blood before it is used in transfusions. The obvious benefit of these tests is the reduced incidence of infection through blood transfusions. The report indicates that the current price of decontaminated blood is $60 per pint. However, if the new screening methods are adopted, the demand and supply for decontaminated blood will change to Q^d= 210 - 1.5P and Q^s = 2.5P - 150. What do you expect to prevail if the new screening methods are adopted? How many units of blood will be used in the United States? What is the level of consumer and producer surplus? Illustrate your findings in a graph.
Answer #1

Answer The distributes to hospitals so stales recent report indicales nearly so Americans contact de contam blood will change= 225-150 = 75 unil a la The consumer Surplus the amount of proful Consumer paying low price commodity williger while prducedProduces are ready sell commodity below to budget ao for Same Commodity surplus of producers Consumers for the commodi The erProduces surplas Amount = 1 base xheigh = ļ 75 + (90-60) = 4x75 (30) I x 2250 1,125